# On the correlation between Architectural Smells and Static Analysis Warnings

### v1

@@ 8 Conclusion
Our results show a statistically significant positive correlation between SAWs and AS, showing weak to moderate correlations across different tools and SAWs and AS in 661 out of 663 tests. In terms of co-occurrence, we found a statistically significant difference in the pairs of SAWs and ASs, with 88% of the Wilcoxon All-Pair Ranking Test cases rejecting the null hypothesis. Specifically, we showed that SAWs can influence the presence of ASs. When prioritizing architectural smells, our findings indicate that using a prioritization method based on the probability of SAW occurrences ($\mathcal{P}$) can effectively guide remediation efforts, revealing a statistically significant difference between the three ranking methods tested. Our proposed approach to SAW prioritization highlights that, in the absence of AS indication, SAW severity and our empirically computed $\mathcal{P}$ together with the possibility of dropping a third of the data to inspect, proved to be effective ASs remediation effort prioritization.
Future research efforts should enhance SAT capabilities to detect the most AS-prone SAWs more accurately, improving AS remediation prioritization. Additionally, research should aim to deepen the analysis of SAWs’ AS-proneness to better trace AS back to SAWs. This will help understand the motivations or root causes linking specific SAWs to specific AS.

### v2

@@ 8 Conclusion
Our results show a statistically significant positive correlation between warnings and smells, showing weak to moderate correlations across different tools and warnings and smells in 661 out of 663 tests. In terms of co-occurrence, we found a statistically significant difference in the pairs of warnings and smells, with 88% of the Wilcoxon All-Pair Ranking Test cases rejecting the null hypothesis. Specifically, we showed that warnings can influence the presence of smells. When prioritizing architectural smells, our findings indicate that using a prioritization method based on the probability of warning occurrences ($\mathcal{P}$) can effectively guide remediation efforts, revealing a statistically significant difference between the three ranking methods tested. Our proposed approach to warning prioritization highlights that, in the absence of smell indications, warning severity and our empirically computed $\mathcal{P}$ together with the possibility of dropping a third of the data to inspect, proved to be an effective smells remediation effort prioritization.
Future research efforts should enhance SAT capabilities to detect the most smell-prone warnings more accurately, improving smell remediation prioritization, and expanding the smells detectable by ARCAN. Additionally, research should aim to deepen the analysis of warnings’ smell-proneness to better trace smells back to warnings. This will help understand the motivations or root causes linking specific warnings to specific smells.
Declarations
Author Contributions: We specify our contributions according to the CRediT taxonomy: Matteo Esposito: Methodology, Writing – Original Draft Preparation. Matteo developed the primary research methodology and contributed significantly to writing the initial manuscript draft. Mikel Robredo: Writing – Original Draft Preparation, Data Curation. Mikel provided essential data preparation and contributed to drafting specific sections. Francesca Arcelli Fontana: Validation, Writing – Review & Editing, Supervision. Francesca critically validated the study’s results, supervised the research’s methodological approach, and reviewed the manuscript for academic rigor. Valentina Lenarduzzi: Conceptualization, Methodology, Supervision, Validation, Writing – Review & Editing. Valentina led the conceptualization and methodological framing of the research, supervised the entire project, validated findings, and contributed to reviewing and refining the manuscript.
Competing Interest: We declare that we have no competing interests.
Data Availability Statement: We provide a replication package on Zenodo [^8] https://zenodo.org/doi/10.5281/zenodo.11366846, which includes the raw data, comprehensive instructions for executing the SAT, a complete list of projects identified as containing warnings and smells, as well as the full results of the statistical analyses, covering both the normality assessments and the hypothesis tests related to our research questions.
Funding: This research received no external funding and was supported solely by the authors’ institutions.